# Reputation in the Shadow of Exit

### v4

@@ 2.2 The Repeated Game
The long-run player plays the stage game in each of infinitely many periods ($t=0,1,2,\dots$), discounting the future at some constant rate $\delta\in(0,1)$. In each period, they face a population of short-run players who take myopic best replies. Short-run player actions are monitored according to $y\in\rho(\cdot|a)\in\Delta(Y)$ for a finite set [^7] Short-run monitoring is irrelevant for the analysis and included only for generality., while long-run player actions are perfectly monitored. Time-$t$ histories are elements $h^{t}=(y_{s},c_{s})_{s=0}^{t-1}$ of past short-run player (and long-run player) actions. Let $H^{t}$ denote the set of time-$t$ histories, $H=\bigcup_{t}H_{t}$ the set of all histories, and $H^{\infty}=(Y\times C)^{\infty}$ the set of all infinite histories. Short-run player strategies are functions $\sigma_{i}:H\to\Delta(A_{i})$, while the long-run player’s strategy is a function $\sigma_{L}:H\to\Delta(C)$ [^8] Following Mertens et al. (2015), there is a canonical mapping from strategies $\sigma:H\to\Delta(\mathcal{X})$ and strategies $\sigma:H^{\infty}\to\Delta(\mathcal{X})$ for $\mathcal{X}\in\{C,A\}$. When it is useful (i.e. when topologizing $\sigma$), I will use these two codomains interchangeably..
@@
$\mathcal{V}(\{a_{t},c_{t}\})=\sum_{t=0}^{\infty}\left(\delta^{t}v(a_{t},c_{t})\prod_{s=0}^{t}f(a_{s},c_{s})\right).$
@@
$V_{*}(\delta)=\sup_{c\in C}\inf_{a\in\mathcal{N}(c)}\sum_{t=0}^{\infty}f(a,c)^{t+1}\delta^{t}v(a,c)$
and their upper Stackelberg strategy as
$V^{*}(\delta)=\sup_{c\in C}\max_{a\in\mathcal{N}(c)}\sum_{t=0}^{\infty}f(a,c)^{t+1}\delta^{t}v(a,c).$
Importantly, we allow the long-run player’s Stackelberg payoffs (and with it, their optimal commitment action) to change with their discount rate. This is because different commitment actions not only affect stage-game payoffs, but also the probability of reaching each terminal node, and hence more patient agents may choose different commitment actions. This feature is special to games with exit and is absent in traditional repeated games models.
Note that the probability of exit in each period is cumulative in the long-run player’s payoff: once they leave the game, their payoff is zero in all future periods regardless of any hypothetical actions that would have been taken. This distinction is a important consequence of exit in my model; in particular, it implies that even as $\delta\to 1$, actions taken in the first few periods still loom large and can have first order impact.
The long-run player can either be a rational type, $\omega_{R}\in\Omega$, who seeks to maximize their discounted payoff, or one of potentially many committed types, $\omega_{c}\in\Omega$, who plays $c$ at every possible history. Short-run players hold a full-support prior belief $\mu_{0}\in\Delta(\Omega)$ over the long-run player’s type.
For any strategy $\sigma$, define an associated belief system $\mu:H\to\Delta(\Omega)$ which outputs some belief about the long-run players’ type for all histories. Together, $(\sigma,\mu)$ is an assessment. Given an assessment, let $\bar{\sigma}_{L}(h^{t})$ denote the unconditional strategy of the long-run player averaging over $\mu(h^{t})$, and $\sigma_{L}(h^{t};\omega)$ denote the strategy of a type $\omega$ long-run player. Analogously, let $\overline{\mathbb{P}}^{\sigma}$ denote the probability measure over histories generated by the unconditional strategy $\sigma$, and $\mathbb{P}^{\sigma}$ the measure generated when $\omega=\omega_{R}$. The main results of this paper study the set of (rational) long-run player payoffs that can be attained in sequential equilibria $(\sigma,\mu)$, defined as follows:
@@
    Short-Run Best Replies. At every history $h^{t}$, $\{\sigma_{i}(h^{t})\}_{i\in[N]}\in\mathcal{N}(\bar{\sigma}_{L}(h^{t}))$.
@@
Because $\sigma:H\to\Delta(C)$ maps an infinite set into another, the notion of convergence is not immediately clear. Say $(\sigma^{n},\mu^{n})$ converges to $(\sigma,\mu)$ if $\sigma^{n}\to\sigma$ converges in the strong topology (see Lemma 3 of Luo and Wolitzky (2025)) and $\mu^{n}\to\mu$ converges weakly. At any history $h^{t}$ where $\mu(h^{t})$ is degenerate, sequential equilibria are equivalent to subgame perfection. Consistency of equilibrium profiles will be useful when evaluating the value of deviations the long-run player may consider which are off the path of play (and not supported by some commitment type); absent some consistency-type condition, the value of such deviations would be hard to define in a principled way. Note the “trembles” defined here for consistency do not change the prior belief over $\Omega$ (which would then change $\mu^{n}$). A standard compactness argument shows this is without loss of generality (I could add the requirement $\mu_{0}^{n}\to\mu_{0}$ as well), but for simplicity I omit this additional notation.
@@
Markov perfection requires the stronger assumption that at any two histories where short-run player posterior beliefs about long-run play are the same, players take the same actions [^9] It may seem unnatural to require the long-run player take the same action when short-run player beliefs are the same. The definition of Markov perfection can be relaxed to require only that short-run players strategies are measurable in $\mu$; however, since there is a single long-run player, their strategy can then be taken to be (without loss of generality) also measurable in $\mu$, less a public randomization device.. While such a refinement may seem a-priori restrictive, I show it arises naturally when studying the effect exit has on reputational incentives; see Theorem 1.
@@
For a fixed discount rate $\delta$, let $\underline{U}(\delta)$ and $\bar{U}(\delta)$ be the long-run player’s lowest and highest payoff from any sequential equilibrium in the game. [^10] There is some potential ambiguity about whether sequential equilibrium need to exist when $C$ is infinite (see Myerson and Reny (2020) for some discussion and general fixes). However, given $\Omega$ is finite and $C$ is countable, the problem is substantially simpler—I show in Proposition 2 and Theorem 1 that sequential equilibria as defined above exist, so in particular $\underline{U},\bar{U}$ are finite. Similarly, let $\underline{U}^{M}(\delta)$ and $\bar{U}^{M}(\delta)$ be the long-run player’s lowest and highest payoff from any MPE in the game.
2.3 Non-normalized payoffs
It will be useful to discuss in somewhat more detail the effect of non-normalized payoff streams on players discounted payoffs, and to highlight why normalization is infeasible as a way to computing average payoffs. Intuitively, without exit, the stream of normalized discounted payoffs converges to the average payoff as $\delta\to 1$. However, when the game has endogenous exit, so long as exit occurs with high enough probability, the average stream of payoffs will be zero for “most” periods (i.e. if the expected exit time is finite), so normalized payoffs will vanish as players get arbitrarily patient. The lemma formalizes this observation.
Lemma 1.
$\lim\limits_{\delta\to 1}(1-\delta)\sum_{t=0}^{\infty}\delta^{t}v_{t}\prod_{s=0}^{\infty}f_{s}=0.$
for any feasible payoffs $v_{t}$.
The proof can be found in the Proof of Lemma 1 Lemma 1 implies that it is not fruitful to compare discounted average payoffs in the patient limit, as they will often be zero even when payoffs are meaningfully different.
This choice, however, leads itself to a novel technical challenge: without normalizing payoffs by the discount rate, deviations today may remain first order (in particular, they are not order $(1-\delta)$ compared to an order $\delta$ variation in the continuation payoff, and are instead order $1$ versus order $\delta$ variation in the continuation payoff). This makes it much harder to discipline the long-run player from deviating to play non-Stackelberg actions, since deviations from potential equilibrium profiles will always remain tempting. Theorem 1 gives a simple economic condition under which the long-run player can successfully establish a (Markovian) reputation, despite this hurdle.
3 Commitment Postures
Let’s start with a basic question intrinsic to our model with exit: How does the discount rate affect the long-run player’s Stackelberg payoff? Define
$c^{*}(\delta)=\operatorname*{\arg\!\max}_{c\in C}\min_{a\in\mathcal{N}(c)}\sum_{t=0}^{\infty}\delta^{t}f(a,c)^{t+1}v(a,c)$
to be the set of (lower) Stackelberg actions for the long-run player when they discount the future at rate $\delta$. Note the long-run player trades off between a higher survival probability at the (potential) cost of a lower stage-game payoff in each period. My first result characterizes exactly how the long-run player resolves this tradeoff as they get more patient: regardless of the speed at which $v(a,c)$ might decline in $c$, they will always favor more aggressive actions that increase their survival probability as they get more patient.
Proposition 1 (Patience Effect).
$c^{*}(\delta)$ is nonempty, upper hemicontinuous, and increasing (in the strong set order) in $\delta$.
Proof.
We first claim $f^{*}(c)$ is continuous, since for any sequence $c_{n}\to c$,
$\lim\limits_{n\to\infty}f^{*}(c_{n})=\lim\limits_{n\to\infty}\frac{v(a^{*}(c_{n}),c_{n})f(a^{*}(c_{n}),c_{n})}{1-\delta f(a^{*}(c_{n}),c_{n})}=\frac{v(a^{*}(c),c)f(a^{*}(c),c)}{1-\delta f(a^{*}(c),c)}.$
The first inequality is by definition, and the second by upper hemi-continuity of the Nash correspondence. Given continuity of $f^{*}(c)$, Berge’s theorem of the maximum applied to the optimizing set $c^{*}(\delta)=\max_{c\in C}f^{*}(c)$ implies $c^{*}(\delta)$ exists and is upper hemi-continuous. Monotonicity now relies on the following lemma, whose proof is deferred to the Proof of Lemma 2  [^11] Such a lemma appears elementary, but to the best of my knowledge there is not a prior reference. Note the lemma establishes a general comparative statics result for discounted streams of payoffs that cross zero once from below.
Lemma 2.
Fix any $\delta$ and $c^{\prime}>c$; if $f^{*}(c)=0$ for all $c\in C$ then the result is trivial; otherwise the solution to $c^{*}(\delta)$ will always lie in $\{c:f^{*}(c)=0\}$. Thus, it is without loss to consider $(c^{\prime},c)$ where $f^{*}(c^{\prime}),f^{*}(c)>0$. Consider the difference
$\sum_{t=0}^{\infty}\delta^{t}[v(a^{*}(c^{\prime}),c^{\prime})f^{*}(c^{\prime})^{t+1}-v(a^{*}(c),c)f^{*}(c)^{t+1}]$
then
The left hand side is strictly positive, while the right hand side is positive, in $(0,1)$ (by Assumption 1), and strictly decreasing in $t$. Thus, there exists a unique $T$ (possibly 0) at which the inequality holds for all $t\geq T$. Lemma 2 then implies, for any $c^{\prime}>c$, $\sum_{t=0}^{\infty}\delta^{t}b_{t}$ is increasing in $\delta$, and hence the objective function $\sum_{t=0}^{\infty}\delta^{t}v(a^{*}(c),c)$ is single-crossing in $(\delta,c)$. The result then follows by the Milgrom and Shannon (1994) comparative statics theorem. ∎
Proposition 1 applies regardless of the speed or curvature of $v$ (in particular, it holds even if $v$ declines very fast relative to $c$), though this comes at the cost of a weak comparative static (every selection from $c^{*}(\delta)$ is nondecreasing in $c$). Note Assumption 2 is not used in the proof. Moreover, it has economic implications in the context of the applications—a regime seeking to fend off a coordinated attack will invest more in survival when they are more patient; longer-lived firms care more about ensures entrants are deterred; patient motorcyclists have a greater incentive to take “separating” actions that prove they are not commitment types, in order to deter exit.
That $c^{*}(\delta)$ varies with the discount rate produces a new order of limits challenge in trying to prove some version of a commitment payoff theorem. In particular, since the long-run player’s Stackelberg action depends on the discount rate, it is not clear that there is a finite set $\Omega$ such that $\Omega\cap c^{*}(\delta)$ is nonempty for all $\delta$ close to $1$. Consequently, for any fixed set of commitment types, it may be infeasible to approximate a patient committed sender.
To obviate this problem, I index and simplify the set of commitment postures, $\Omega(\delta)=\{\omega_{R},\omega_{c^{*}(\delta)}\}$ (where I abuse notation and use $c^{*}(\delta)$ to refer to a single selection from the set of commitment types). Moreover, as before, I set $\underline{U}(\delta)$ and $\bar{U}(\delta)$ as the infimum and supremum of the long-run player’s best and worst sequential equilibrium payoff, though now I implicitly fix the set of commitment types $\Omega=\Omega(\delta)$, though I still suppress dependence on the specific sequence of prior beliefs $\mu_{0}(\delta)\in\Delta(\Omega(\delta))$ (which itself my also change in $\delta$). Give this notation, the limit infimum (supremum) $\liminf_{\delta\to 1}\underline{U}(\delta)$ (resp. $\limsup_{\delta\to 1}$) should be interpreted as the limit infimum (supremum) of the sequence of lower equilibrium payoffs in the repeated game with discount rate $\delta$, commitment types $\Omega(\delta)$, and some fixed sequence of prior beliefs $\mu_{0}(\delta)\in\Delta(\Omega(\delta))^{o}$. This interpretation implies that statements of the form $\limsup_{\delta\to 1}|V_{*}(\delta)-\underline{U}(\delta)|=0$ can be read as “The long-run player attains almost their Stackelberg payoff so long as there is positive probability players believe they are committed to their Stackelberg action”.
@@ 4.1 Complete Information
Consider first the complete information game where there is no uncertainty about the long-run player’s rationality. Recall that a folk theorem holds in the game if, as $\delta\to 1$ the set of payoffs attainable as part of some equilibrium in the complete-information game with discount rate $\delta$ converges to the set of feasible and individually rational payoffs for the long-run player (here the set $V_{0}(\delta),V^{*}(\delta)]$), supposing short-run players’ myopically best reply. Proposition 2 characterizes exactly when a folk theorem holds in our game. Arguments from Fudenberg and Levine (1994), Fudenberg et al. (1994) (which would immediately imply a folk theorem) do not necessarily apply because exit can “loom large,” and thus lead to periods where the effective discount rate is small even as $\delta\to 1$. Thus, there exist games where not all feasible, individually rational payoffs are attainable. The next result characterizes, as a function of long-run player payoffs, exactly when a folk-like theorem holds as a function of the long-run player’s payoff, and formalizes the sense in which efficient payoffs can be sustained only when the future “looms large” given the exit structure. Throughout, let $L(a,c)=\frac{f(a,c)}{f(a,0)}$ be the likelihood ratio of exit when the long-run player chooses $c$ at some action instead of 0 given short-run player action $a$.
@@
Let $\mu_{0}(\omega_{R})=1$ and let $(a^{*},c^{*})$ attain $V^{*}(\delta)$. Then
$\lim\limits_{\delta\to 1}[\underline{U}(\delta),\bar{U}(\delta)]=\lim\limits_{\delta\to 1}[\underline{V}_{0}(\delta),V^{*}(\delta)]$
if and only if there exists $\bar{\delta}<1$ such that for all $\delta>\bar{\delta}$ and any $(a^{*},c^{*})$ attaining $V^{*}(\delta)$,
$\delta[V^{*}(\delta)L(a^{*},c^{*})-\underline{V}_{0}(\delta)]\geq v(a^{*},0)-v(a^{*},c^{*})L(a^{*},c^{*}).$
If either of the two (equivalent) statements above holds, say the game admits a folk theorem. Proposition 2 gives a good baseline for when all feasible payoffs can be obtained in some equilibrium; moreover, the relevant incentive compatibility conditions will give, under some monotonicity and uniqueness conditions, an exact characterization of when the long-run player exactly secures their Stackelberg payoff. Moreover, the converse direction of the result implies the long-run player’s best equilibrium payoff is bounded away from their Stackelberg payoff even as players get arbitrarily patient (i.e. the folk theorem fails) so long as the incentive compatibility condition fails.
The proof is relatively straightforward and can be found in the Proof of Proposition 2 The key step in the proof is to write the incentive constraint for the long-run player, supposing they obtain $V^{*}(\delta)$ if they play $c^{*}$ and $\underline{V}_{0}(\delta)$ otherwise and algebraically rearrange terms to conclude it is equivalent to the condition in the proposition. This then implies playing $c^{*}(\delta)$ is self-enforcing if (and only if) the condition in the result is satisfied.
Importantly, though, the equilibria constructed in the proof of Proposition 2 to sustain a high payoff for the long-run player require that punishments off the equilibrium path are non-Markovian, in the sense that they are not measurable in $\mu_{t}(h^{t})$ (which is constant in the complete information case). Given this observation, it is perhaps unsurprisingly that the Markovian restriction can strongly refine the set of payoffs: in the complete information game, it does so exactly by holding the long-run player down to their exact static payoff.
@@
Fix any $\delta$. Then $\underline{U}^{M}(\delta),\bar{U}^{M}(\delta)]=[\underline{V}_{0}(\delta),\bar{V}(\delta)]$.
This property of the complete-information Markov game will be important in disciplining the value of deviations in the game with reputational concerns. The proof follows immediately once it becomes clear that the Markovian refinement requires the continuation strategies to be independent of the history (since $\mu(h^{t})=1$ at all histories in all sequential equilibria). For completeness, I give a formal argument in the Proof of Proposition 3 . Note that Proposition 3 applies to any repeated game, not necessarily just those with exit.

### v5

@@ 2.2 The Repeated Game
The long-run player plays the stage game in each of infinitely many periods ($t=0,1,2,\dots$), weighting each period equally. In each period, they face a population of short-run players who take myopic best replies.
Time-$t$ histories are elements $h^{t}=(y_{s},c_{s})_{s=0}^{t-1}$ of past short-run player signals and long-run player actions. Let $H^{t}$ denote the set of time-$t$ histories, $H=\bigcup_{t}H_{t}$ the set of all histories, and $H^{\infty}=(Y\times C)^{\infty}$ the set of all infinite histories. Short-run player strategies are functions $\sigma_{i}:H\to\Delta(A_{i})$, while the (rational) long-run player’s strategy is a function $\sigma_{L}:H\to\Delta(C)$. [^8] Following Mertens et al. (2015), there is a canonical mapping from strategies $\sigma:H\to\Delta(\mathcal{X})$ and strategies $\sigma:H^{\infty}\to\Delta(\mathcal{X})$ for $\mathcal{X}\in\{C,A\}$. When it is useful (i.e. when topologizing $\sigma$), I will use these two codomains interchangeably.
@@
$\mathcal{V}(\{a_{t},c_{t}\})=\sum_{t=0}^{\infty}\left(v(a_{t},c_{t})\prod_{s=0}^{t}f(a_{s},c_{s})\right).$
Note payoffs here are not discounted. SUPPLEMENTAL APPENDIX studies the discounted repeated game and shows the game here is the standard limit as players become patient, but with exit. However, discounting payoffs does not add anything to the analysis and, given the exit structure, would in fact substantially complicate the analysis. Consequently, I omit the discount rate $\delta$ in the main body of the paper, and refer to this as the case of “extreme patience” for the long-run player.
@@
$V_{*}=\sup_{c\in C}\min_{a\in\mathcal{N}(c)}\sum_{t=0}^{\infty}f(a,c)^{t+1}v(a,c)=\sup_{c\in C}\min_{a\in\mathcal{N}(c)}\frac{f(a,c)v(a,c)}{1-f(a,c)}$
and let $c^{*}$ be a maximizer in $V_{*}$ (throughout, suppose such a maximizer exists). Let $V^{*}$ be the upper Stackelberg payoff, defined analogously.
Payoffs are written such that exit is permanent: once exit has occurred, the long-run player leaves the game, and hence their payoff is zero in all future periods regardless of any hypothetical actions that would have been taken.
The long-run player can either be a rational type, $\omega_{R}\in\Omega$, who seeks to maximize their discounted payoff, or one of potentially many committed types, $\omega_{c}\in\Omega$, who plays $c$ at every possible history. Short-run players hold a full-support prior belief $\mu_{0}\in\Delta(\Omega)$ over the long-run player’s type. Write $\sigma(\cdot,\omega_{c})\equiv c$ for the strategy of a commited type $\omega_{c}$.
For any strategy $\sigma$, define an associated belief system $\mu:H\to\Delta(\Omega)$ which outputs some belief about the long-run players’ type for all histories. Together, $(\sigma,\mu)$ is an assessment. Given an assessment, let $\overline{\sigma}_{L}(h^{t})$ denote the unconditional strategy of the long-run player averaging over $\mu(h^{t})$, and $\sigma_{L}(h^{t};\omega)$ denote the strategy of a type $\omega$ long-run player. Analogously, let $\overline{\mathbb{P}}^{\sigma}$ denote the probability measure over histories generated by the unconditional strategy $\sigma$, and $\mathbb{P}^{\sigma}$ the measure generated when $\omega=\omega_{R}$. The main results of this paper study the set of (rational) long-run player payoffs that can be attained in sequential equilibria $(\sigma,\mu)$, defined as follows:
@@
    Short-Run Best Replies. At every history $h^{t}$, $\{\sigma_{i}(h^{t})\}_{i\in[N]}\in\mathcal{N}(\overline{\sigma}_{L}(h^{t}))$.
@@
Because $\sigma:H\to\Delta(C)$ maps an infinite set into another, the notion of convergence is not immediately clear. Say $(\sigma^{n},\mu^{n})$ converges to $(\sigma,\mu)$ if $\sigma^{n}\to\sigma$ converges in the strong topology (see Appendix B of Luo (2025)) and $\mu^{n}\to\mu$ converges weakly. At any history $h^{t}$ where $\mu(h^{t})$ is degenerate, sequential equilibria are equivalent to subgame perfect equilibria. Consistency of equilibrium profiles will be useful when evaluating the value of deviations the long-run player may consider which are off the path of play (and not supported by some commitment type); absent some consistency-type condition, the value of such deviations would be hard to define in a principled way. Note the “trembles” defined here for consistency do not change the prior belief over $\Omega$ (which would then change $\mu^{n}$). A standard compactness argument shows this is without loss of generality (one could add the requirement $\mu_{0}^{n}\to\mu_{0}$ as well), but for simplicity I omit this additional notation.
@@
Markov perfection requires the stronger assumption that at any two histories where short-run player posterior beliefs about long-run play are the same, players take the same actions. [^9] It may seem unnatural to require the long-run player take the same action when short-run player beliefs are the same. The definition of Markov perfection can be relaxed to require only that short-run players strategies are measurable in $\mu$; however, since there is a single long-run player, their strategy can then be taken to be (without loss of generality) also measurable in $\mu$, less a public randomization device. While such a refinement may seem a-priori restrictive, I show it arises naturally when studying the effect exit has on reputational incentives (see Theorem 1).
@@
Let $\underline{U}$ and $\overline{U}$ be the long-run player’s lowest and highest payoff from any sequential equilibrium in the game. [^10] There is some potential ambiguity about whether sequential equilibrium need to exist when $C$ is infinite (see Myerson and Reny (2020) for some discussion and general fixes). However, given $\Omega$ is finite and $C$ is countable, the problem is substantially simpler—I show in Proposition 1 and Theorem 1 that sequential equilibria as defined above exist, so in particular $\underline{U},\overline{U}$ are finite. Similarly, let $\underline{U}^{M}$ and $\overline{U}^{M}$ be the long-run player’s lowest and highest payoff from any MPE in the game.
@@ 3 Commitment Payoff Theorems
This section contains the main results of this paper, which characterize the payoff set in the complete information game with and without reputation, in sequential and Markovian equilibria. Informally, I establish the following table, which summarizes the takeaway from the results in this section. Together, these results form the main contribution of this paper.
non-Markov Markov
Reputation Anti-Commitment Theorem (Prop. 3) Stackelberg Payoff (Theorem 1)
No Reputation Folk Theorem (Prop. 1) Static Nash Payoff (Prop. 2)
@@ 3.1 Complete Information
Consider the complete information game where there is no uncertainty about the long-run player’s rationality. Say a folk theorem holds in the game if the set of payoffs attainable as part of some equilibrium in the complete-information game is the set of feasible and individually rational payoffs for the long-run player (here the set $V_{0},V^{*}]$), supposing short-run players’ myopically best reply. Proposition 1 characterizes exactly when a folk theorem holds in our game. Arguments from Fudenberg and Levine (1994), Fudenberg et al. (1994) (which would immediately imply a folk theorem) do not necessarily apply because exit can “loom large,” and thus lead to periods where the effective discount rate is small even without discounting, i.e. $\delta=1$. Thus, there exist games where not all feasible, individually rational payoffs are attainable. Proposition 1 characterizes, as a function of long-run player payoffs, exactly when a folk-like theorem holds as a function of the long-run player’s payoff, and formalizes the sense in which efficient payoffs can be sustained only when the future “looms large” given the exit structure. Let $L(a,c)=\frac{f(a,c)}{f(a,0)}$ be the likelihood ratio of exit when the long-run player chooses $c$ at some action instead of 0 given short-run player action $a$.
@@
Let $\mu_{0}(\omega_{R})=1$ and let $(a^{*},c^{*})$ attain $V^{*}$. Then
$\underline{U},\overline{U}]=[\underline{V}_{0},V^{*}]$
if and only if for any $(a^{*},c^{*})$ attaining $V^{*}$,
$(V^{*})L(a^{*},c^{*})-\underline{V}_{0}\geq v(a^{*},0)-v(a^{*},c^{*})L(a^{*},c^{*}).$
If either of the two (equivalent) statements above holds, say the game admits a folk theorem. Proposition 1 gives a good baseline for when all feasible payoffs can be obtained in some equilibrium; moreover, the relevant incentive compatibility in Proposition 1 will give, under some monotonicity and uniqueness conditions, an exact characterization of when the long-run player exactly secures their Stackelberg payoff (Corollary 1). The converse direction of the result implies the long-run player’s best equilibrium payoff is bounded away from their Stackelberg payoff even as players get extremely patient (i.e. the folk theorem fails) so long as the incentive compatibility condition fails.
The proof is relatively straightforward and can be found in the PROOF OF PROPOSITION 1 The key step in the proof is to write the incentive constraint for the long-run player, supposing they obtain $V^{*}$ if they play $c^{*}$ and $\underline{V}_{0}$ otherwise and algebraically rearrange terms to conclude it is equivalent to the condition in the proposition. This then implies playing $c^{*}$ is self-enforcing if (and only if) the condition in the result is satisfied.
Importantly, though, the equilibria constructed in the proof of Proposition 1 to sustain a high payoff for the long-run player require that punishments off the equilibrium path are non-Markov, in the sense that they are not measurable in $\mu_{t}(h^{t})$ (which is constant in the complete information case). Given this observation, it is perhaps unsurprisingly that the Markov restriction can strongly refine the set of payoffs: in the complete information game, it does so exactly by holding the long-run player down to their exact static payoff.
@@
Let $\mu_{0}(\omega_{R})=1$. Then $\underline{U}^{M},\overline{U}^{M}]=[\underline{V}_{0},\overline{V}_{0}]$.
This property of the complete-information Markov game will be important in disciplining the value of deviations in the game with reputational concerns. The proof follows immediately once it becomes clear that the Markov refinement requires the continuation strategies to be independent of the history (since $\mu(h^{t})=1$ at all histories in all sequential equilibria). For completeness, I give a formal argument in the PROOF OF PROPOSITION 2 . Note that Proposition 2 applies to any repeated game, not necessarily just those with exit.